A diffraction grating is made up of slits of width a with separation d. The grating is illuminated by monochromatic plane waves of wavelength λ at normal incidence. What is the angular width of a spectral line observed in the first order if the grating has N slits? State your answer in terms of the given variables.
Answer #1

slit width = a

slits separation = d

wavelength = λ

angular width = ?

We know that    d*sin(θ) = m*λ

angular width = sin(θ) = m*λ/d

For first order, m = 1

Then angular width = sin(θ) = λ/d